OpenStudy (daniel.ohearn1):

The Integral of (ln(x))^x dx seems easy enough but to express it indefinitely?

OpenStudy (pawanyadav):

Isn't it ln(x^x)dx ...

ganeshie8 (ganeshie8):

could you provide some context of this problem ?

OpenStudy (kainui):

Might just be a coincidence, but I noticed these graphs seem to be approximately the same: \[(\ln x)^x \approx e^{x-e}\] https://www.desmos.com/calculator/dsxiahpiw6 Of course one of these is much easier to integrate than the other, so it would be nice if there was some definite relationship between the two.

Miracrown (miracrown):

Shouldn't it be in(x^x)dx @daniel.ohearn1

Miracrown (miracrown):

We can use a property of logarithms xln(x)dx we pull down the exponent

Miracrown (miracrown):

any ideas about how to integrate xln(x) ?

Miracrown (miracrown):

so we need to assign u and dv... what could we let u be ?

OpenStudy (bobo-i-bo):

If it is (lnx)^x, then note that (lnx)^x=e^(x*ln(lnx))

Miracrown (miracrown):

Yes, I agree with that.

OpenStudy (agent0smith):

Something tells me the integral of (ln x)^x is not "easy enough": http://m.wolframalpha.com/input/?i=integral+of+%28ln+%28x%29%29%5Ex&x=0&y=0

OpenStudy (daniel.ohearn1):

No it isn't ln(x^x) nor is it "impossible". But it could be unique, and there could be a umbrella function for it..

OpenStudy (bobo-i-bo):

It's convincing enough for me :P